OpenStudy (anonymous):

differential to estimate the value (26)^(2/3)

OpenStudy (anonymous):

The answer is 8.78. I used my calculator.

OpenStudy (anonymous):

i guess you are supposed to use \(27\) since \(27^{\frac{2}{3}}=9\)

OpenStudy (anonymous):

idea is to linearize \[\sqrt[3]{(27-x)^2}\] and then replace \(x\) by 1

OpenStudy (anonymous):

Use differentials to estimate the value of the indicated expression. Then compare your estimate with the result given by a calculator.

OpenStudy (anonymous):

take the derivative, replace \(x\) by 27 to find the slope, then use the point slope formula to find the equation of the line tangent to the graph of \(y=\sqrt[3](27-x)^2}\) at the point \((27,9)\)

OpenStudy (anonymous):

that should be \(y=\sqrt[3]{27-x)^2} \)